2. Radiation of a certain wavelength causes electrons with a maximum kinetic energy of 0.66 eV to be ejected from a metal whose work function is 2.50 eV. What will be the maximum kinetic energy with which this same radiation ejects electrons from another metal whose work function is 2.23 eV?
eV
Answer #1

According to Einstein hv = work function + max. Kinetic energy

Where hv energy of incident wave  

hv= 2.5 + 0.66 = 3.16eV

So for work function 2.23eV

3.16 = 2.23 + max. Kinetic energy

Max. Kinetic energy = 3.16 - 2.23 = 0.93 eV